What is the reference angle in radians for -3Pi/8
Answer by stanbon(75887)

What is the reference angle in radians for -3Pi/8
---
The reference angle is the acute angle made by
the terminal side of the given angle and the
closest part of the x axis.
----------------------------------
Your problem:
reference = 0-(-(3/8)pi) = (3/8)pi
==================================
Cheers,
Stan H.
=================